A sperical balloon of 21cm diameter is to be filled with hydrogen at NTP from a cylinder containing the gas at 20 atm and 27°C. if the cylinder can hold 2. 82L of water, calculate the number of balloons tat can be filled up..

Grade:12th pass

1 Answers

Umakant biswal
7 years ago
@ sowmya 
here i am giving u the steps which u can apply and solve the questions 
see, here we will know the volume of the baloon . and the pressure inside it , we will hv to assume it to be 1 atm ., beacuse if more then the baloon will explode . 
so, now we can find the mass of the air in the baloon 
by applying the ideal gas equation 
PV = NRT 
N = m/M 
SO, PV = m/M *RT 
m= PVM /RT 
LET m for the baloon m1 . 
now we got the mass of hydrogen in each baloon . now, lets move to the tank . 
we know the volume , initial pressure and temp . so, we can find the initial mass of hydrogen in the tank . 
initial here is 20 atm pressure , rest all same , above formula applied . 
keep volume as 2.82 l and apply the same formula . 
now , obviously , cylinder have to be some gas left in it , it cannot be fully empty . 
when atm , and pres inside cylinder will equal , then gas will not move anymore . 
NOW , U GOT INITIAL AND FINAL MASS OF GAS IN CYLINDER , SUBSTRACT IT . 
YOU WILL GET THE MASS TRANSFERRED TO BALOON 
FINAL MASS IS 1 ATM AND 2.82 L 
CYLINDER WILL HAVE SOME GAS LEFT BEHIND . 
THEN , DIVIDE THE TRANSFERRED MASS BY THE MASS OF AIR , IN EACH BALOON . "
WHICH U CALCULATED IN FRST STEP . "
SO, U HAVE TO CALCULATE 3 MASSES IN TOTAL . 
THE INITIAL MASS IS 20 TIMES MORE OF FINAL MASS . 
i think u can do it now , i am just giving the value below 
p(initial) = 20 atm , p (final ) = 1 atm 
v (initial ) = 8.72l . v (final ) = 8.72 l 
then apply formula,,
m= PVM / RT 
hope it helps 
if confused in any step , let me know 
all the best ..